# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ VI.2.1 An experimental approach
These three temperature cycles are strong evidence for this interpretation of memory. This is at odds with Bouchaud et al. [67], who argue that the coherent volume returns to the native value upon reheating. Adjusting $t_{\mathrm{w2}}$, one can change the value of the coherent volume when returning to $T_{1}$ at will, from memory loss to complete loss of memory. This behavior is born out in Fig. 48, and also in the independent experiments and model of Freedberg et al. [125].
@@ VI.2.2 Rejuvenation and memory in computer experiments
Recall that in native runs the peak of $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ defines $t^{\mathrm{eff}}_{\mathrm{w}}\simeq t_{\mathrm{w}}$ [compare Fig. 35 with Fig. 49 (left)].
@@
where the scaling law is based on Eq. (70). Consider two cases that differ in the cold temperature $T_{2}$: $T_{2}=0.5$ meets the chaos requirement of Eq. (73); $T_{2}=0.7$ does not [recall that the glass temperature is $T_{\mathrm{g}}=1.102(3)$].
The main result of Baity-Jesi et al. [31] is reproduced in Fig. 49. Rejuvenation is well visible in the central-bottom panel, as the peak of $S_{\mathrm{ZFC}}$ occurs at time $t^{\mathrm{eff}}_{\mathrm{w}}\ll(t_{\mathrm{w1}}+t_{\mathrm{w2}})$. This analysis is phenomenological and closely related to the experimental one, but the advantage of “numerical experiments” is that they give access to the full configurations and, hence, to microscopic correlation functions. To explain this phenomenon, we need three length scales: $\xi_{\mathrm{micro}}$, $\zeta$ and $\xi_{\mathrm{Zeeman}}$ (see Sec. I.3). Let us focus on Fig. 50. In a fixed-temperature protocol (native runs) $\xi_{\mathrm{Zeeman}}$ follows the behavior of $\xi_{\mathrm{micro}}$, recall Fig. 32. In contrast, in a jump run, the situation is more intricate. The size of the glassy domains, i.e., $\xi_{\mathrm{micro}}$, is frozen, but the system responds to the external field with a shorter $\xi_{\mathrm{Zeeman}}$ (rejuvenation). In particular, if the cold temperature $T_{2}$ satisfies Eq. (73), we have full chaos and $\xi_{\mathrm{Zeeman}}^{\mathrm{jump}}(T_{2},t_{\mathrm{w2}})\sim\xi_{\mathrm{Zeeman}}^{\mathrm{native}}(T_{2},t_{\mathrm{w2}}-t_{\mathrm{w1}})$; otherwise, we have an intermediate situation, as the $T_{2}=0.7$ case shows. The last ingredient is that $\xi_{\mathrm{Zeeman}}^{\mathrm{jump}}$ values can be matched by a simple factor to the values of $\zeta^{\mathrm{jump}}$. This implies that, in the presence of strong chaos, the defects are the ones responding to the magnetic field. This observation will be vital for the quantification of memory, as it allows us to perform numerical experiments in absence of the external field and simply calculate $\zeta$ to quantify $\xi_{\mathrm{Zeeman}}$ in temperature-varying protocols.
@@ VI.2.3 Simulation and experiment: the memory coefficients
The equilibrium non-linear susceptibility is proportional to the integral of $r^{2}C_{4}$ for $0<r<\infty$ [63]. Thus, following Baity-Jesi et al. [36, 33], Zhai et al. [302], Paga et al. [232], one can generalize this equilibrium relation by computing these integrals using the non-equilibrium correlation function $C_{4}$ of Eq. (22). As shown in Fig. 52, there is a small but detectable difference in the behavior of $r^{2}C_{4}$ as $t_{\mathrm{w2}}$ varies in temperature cycles. Paga et al. [230] considered the curve with $t_{\mathrm{w2}}=0$ as the reference curve and proposed as numerical memory coefficient
@@ VI.3 A different kind of memory: the Mpemba effect
Two different off-equilibrium systems are simulated. In the first one, the temperature of the thermal reservoir is quenched from infinite temperature to $T_{1}=1.3$, which is above the critical temperature [$T_{\mathrm{g}}=1.102(3)$, see Baity-Jesi et al. [27]]. The system is left to evolve until it reaches the energy $e_{1}\approx-1.6428$ and this is the starting point of the first system in the numerical experiment. The second simulated system is prepared in a very similar way but the temperature of the thermal reservoir is now $T_{2}=1.2$ and the initial energy is lower, $e_{2}=-1.67174$.
@@ VII.1 What is yet to be nailed down theoretically
On the low-temperature phase.
The renormalization group (RG), which has served us so well to clarify the nature of phases and phase transitions in clean systems [291, 77, 14], needs to be extended to spin glasses. This will not be easy. Numerical simulations clearly hint that universality (perhaps the most important consequence of the RG framework) extends to disordered systems [39, 126, 128, 127], including classical spin glasses [145, 107]. For quantum spin glasses we even have direct experimental confirmation [177] of critical exponents obtained with numerical simulations [51].
Furthermore, it is very difficult to determine from experimental and numerical analysis the different universality classes, due to the large uncertainties in the values of the critical exponents. One of the exceptions to this situation was the computation of the critical exponents in the three-dimensional Ising Edwards-Anderson model by Baity-Jesi et al. [27]. As discussed in detail in this review, the Ising Edwards-Anderson model beautifully describes the non-equilibrium dynamics of CuMn samples, even if CuMn is a Heisenberg material. This discrepancy can be reconciled by noticing that small but unavoidable anisotropies in the spin-spin interactions lead to Ising universality [37]. However, not all is understood, since experimental measurements for critical exponents in spin glasses are not close to the Ising values. For example, the non-linear magnetization exponent $\gamma=6.33(11)$ was extracted from simulations of the three-dimensional Ising spin glass [27]. Probably the most accurate experimental determination was obtained from zero-field measurements on the Heisenberg spin glass 0.5 at.AgMn [190] and yielded a very different $\gamma=2.3\pm 0.2$. The authors are unaware of other experimental reports for values of $\gamma$ that come close to the value obtained from simulations. Similar differences between simulation and experiment are present for other critical exponents (e.g., $\beta,~{}\delta,~{}{\text{and}}~{}\nu$).
@@
Some new ideas to carry out non-perturbative computations would clearly be welcome. These might include further development of real-space tools such as the Monte Carlo RG [241] or the $M$-expansion around the Bethe lattice [8, 17]. More speculatively, we could think of the metastate —the mathematically rigorous way of building the low-temperature phase [6, 222, 223, 224, 249], see Newman et al. [221] for a recent review. It is conceivable that the De Dominicis RG program at zero field  [98, 95] could blend with the metastate approach to build a rigorous theory of the spin-glass phase [249].
On numerical simulations.
As discussed in Sec. II.7, NP-completeness is making spin glasses increasingly popular in an engineering context. This has driven the development of specific hardware to solve spin-glass problems. One may easily speculate that the main difficulties hampering spin-glass numerical simulations are common to both the physics and the engineering contexts. It follows that eventual advances in simulation algorithms (or hardware) in physics may have far-reaching consequences.
@@
Beside new hardware developments (see Sec. II.7), new simulation algorithms are needed. Temperature chaos suggests that these should work at a single temperature. Alternatively, ingenious ways of circumventing chaotic phenomena should be devised. Quantum annealing could be such an approach, but it may be hampered by chaotic phenomena related to the transverse field [180].
Experiments with potential for strong impact on theory.

### v2

@@ VI.2.1 An experimental approach
These three temperature cycles are strong evidence for our interpretation of memory. This is more quantitative than Bouchaud et al. [72], who arrive at the same conclusion but only state that “memory is not perfect if $t_{\mathrm{w2}}$ is large enough.” Adjusting $t_{\mathrm{w2}}$, one can change the value of the coherent volume when returning to $T_{1}$ at will, from memory loss to complete loss of memory. This behavior is born out in Fig. 49, and also in the independent experiments and model of Freedberg et al. [147].
@@ VI.2.2 Rejuvenation and memory in computer experiments
Recall that in native runs the peak of $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ defines $t^{\mathrm{eff}}_{\mathrm{w}}\simeq t_{\mathrm{w}}$ [compare Fig. 36 with Fig. 50 (left)]. The numerical relaxation functions also exhibit a spurious peak at $t\sim 2^{10}$, which lacks any physical significance. We believe this peak arises from a spin-coupling effect that occurs as soon as the external magnetic field is applied. This hypothesis is supported by the observation that the short-time peak does not shift with $t_{\mathrm{w}}$. Moreover, for small $t_{\mathrm{w}}\sim 2^{10}$, this spurious peak merges with the physical one, making the extraction of $\xi_{\mathrm{Zeeman}}$ even more challenging —unless one employs the new strategy introduced in Sec. V.3.
@@
where the scaling law is based on Eq. (74). Consider two cases with $T_{1}=0.9$ that differ in the cold temperature $T_{2}$: $T_{2}=0.5$ meets the chaos requirement of Eq. (77); $T_{2}=0.7$ does not [recall that the glass temperature is $T_{\mathrm{g}}=1.102(3)$].
The main result of Baity-Jesi et al. [33] is reproduced in Fig. 50. Rejuvenation is well visible in the central-bottom panel, as the peak of $S_{\mathrm{ZFC}}$ occurs at time $t^{\mathrm{eff}}_{\mathrm{w}}\ll(t_{\mathrm{w1}}+t_{\mathrm{w2}})$. This analysis is phenomenological and closely related to the experimental one, but the advantage of “numerical experiments” is that they give access to the full configurations and, hence, to microscopic correlation functions. To explain this phenomenon, we need three length scales: $\xi_{\mathrm{micro}}$, $\zeta$ and $\xi_{\mathrm{Zeeman}}$ (see Sec. I.3). Let us focus on Fig. 51. In a fixed-temperature protocol (native runs) $\xi_{\mathrm{Zeeman}}$ follows the behavior of $\xi_{\mathrm{micro}}$, recall Fig. 33. In contrast, in a jump run, the situation is more intricate. The size of the glassy domains, i.e., $\xi_{\mathrm{micro}}$, is frozen, but the system responds to the external field with a shorter $\xi_{\mathrm{Zeeman}}$ (rejuvenation). In particular, if the cold temperature $T_{2}$ satisfies Eq. (77), we have full chaos and $\xi_{\mathrm{Zeeman}}^{\mathrm{jump}}(T_{2},t_{\mathrm{w2}})\sim\xi_{\mathrm{Zeeman}}^{\mathrm{native}}(T_{2},t_{\mathrm{w2}}-t_{\mathrm{w1}})$; otherwise, we have an intermediate situation, as the $T_{2}=0.7$ case shows. The last ingredient is that $\xi_{\mathrm{Zeeman}}^{\mathrm{jump}}$ values can be matched by a simple factor to the values of $\zeta^{\mathrm{jump}}$. This implies that, in the presence of strong chaos, the rearrangements are the ones responding to the magnetic field. This observation will be vital for the quantification of memory, as it allows us to perform numerical experiments without an external field and simply calculate $\zeta$ to quantify $\xi_{\mathrm{Zeeman}}$ in temperature-varying protocols.
@@ VI.2.3 Simulation and experiment: the memory coefficients
The equilibrium nonlinear susceptibility is proportional to the integral of $r^{2}C_{4}$ for $0<r<\infty$ [66]. Thus, following Baity-Jesi et al. [38, 35], Zhai et al. [376], Paga et al. [289], one can generalize this equilibrium relation by computing these integrals using the nonequilibrium correlation function $C_{4}$ of Eq. (24). As shown in Fig. 53, there is a small but detectable difference in the behavior of $r^{2}C_{4}$ as $t_{\mathrm{w2}}$ varies in temperature cycles. Paga et al. [287] considered the curve with $t_{\mathrm{w2}}=0$ as the reference curve and proposed as numerical memory coefficient
@@ VI.3 A different kind of memory: the Mpemba effect
Two different off-equilibrium systems are simulated. In the first one, the temperature of the thermal reservoir is quenched from infinite temperature to $T_{1}=1.3$, which is above the critical temperature $T_{\mathrm{g}}=1.102(3)$, see Baity-Jesi et al. [29]]. The system is left to evolve until it reaches the energy $e_{1}\approx-1.6428$ and this is the starting point of the first system in the numerical experiment. The second simulated system is prepared in a very similar way but the temperature of the thermal reservoir is now $T_{2}=1.2$ and the initial energy is lower, $e_{2}=-1.67174$.
@@ VII.1 What is yet to be nailed down theoretically
VII.1.1 On the low-temperature phase.
The renormalization group (RG), which has served us so well to clarify the nature of phases and phase transitions in clean systems [366, 85, 16], needs to be extended to spin glasses. This will not be easy. Numerical simulations clearly hint that universality —perhaps the most important consequence of the RG framework— extends to disordered systems [41, 148, 150, 149], including classical spin glasses [169, 124]. For quantum spin glasses we even have direct experimental confirmation [211] of critical exponents obtained with numerical simulations [53].
@@
Some new ideas to carry out nonperturbative computations would clearly be welcome. These might include further development of real-space tools such as the Monte Carlo RG [299] or the $M$-expansion around the Bethe lattice [10, 19]. More speculatively, we could think of the metastate —the mathematically rigorous way of building the low-temperature phase [6, 277, 278, 279, 313], see Newman et al. [276] for a recent review. It is conceivable that the De Dominicis RG program at zero field  [113, 110] could blend with the metastate approach to build a rigorous theory of the spin-glass phase [313].
VII.1.2 On numerical simulations
As discussed in Sec. II.8, NP-completeness is making spin glasses increasingly popular in an engineering context. This has driven the development of specific hardware to solve spin-glass problems. One may easily speculate that the main difficulties hampering spin-glass numerical simulations are common to both the physics and the engineering contexts. It follows that eventual advances in simulation algorithms (or hardware) in physics may have far-reaching consequences.
@@ VII.1.2 On numerical simulations
Besides new hardware developments (see Sec. II.8), new simulation algorithms are needed. Temperature chaos suggests that these should work at a single temperature. Alternatively, ingenious ways of circumventing chaotic phenomena should be devised. Quantum annealing could be such an approach, but it may be hampered by chaotic phenomena related to the transverse field [214].
VII.1.3 The chiral-glass scenario and the problem of the critical exponents
The alert reader will have noticed an important paradox. On the one hand, as discussed in detail in this review, the Ising Edwards-Anderson model beautifully describes the nonequilibrium dynamics of CuMn samples below $T_{\mathrm{g}}$, even if CuMn is a Heisenberg material (more on this below). On the other hand, the critical exponents of the three-dimensional Ising Edwards-Anderson model [which are quite well known by now, see 29], do not agree with their experimental counterpart. Take, for instance, the nonlinear magnetization exponent $\gamma$. Its value for the Ising Edwards-Anderson model is $\gamma=6.33\pm 0.11$. Probably the most accurate experimental determination of the same critical exponent was obtained from zero-field measurements on the Heisenberg spin glass 0.5 at.% AgMn [227] and yielded a very different $\gamma=2.3\pm 0.2$. The authors are unaware of other experimental reports for values of $\gamma$ in metallic spin glasses that come close to the value obtained from simulations. Similar differences between simulation and experiment are present for other critical exponents (e.g., $\beta,~\delta,~{\text{and}}~\nu$).
The failure in reproducing the critical exponents to describe CuMn or AgMn at temperatures near to, but above, $T_{\mathrm{g}}$ forces one to consider what could be the simplest, yet adequate, model to describe this temperature region. In fact, microscopic models of spin glasses include several interactions, such as the RKKY interaction, which is invariant over rotations [320, 194, 370] and the Dzyaloshinsky-Moriya (DM) interaction that breaks the rotational symmetry [120, 264]. The DM interaction introduces a certain degree of random anisotropy, which results in real spin glasses that are never fully isotropic (this theoretical limit is named the Heisenberg spin glass). It is possible to classify materials according to the degree of anisotropy of their interactions [310]. Perhaps the best example of an Ising spin glass is the extremely anisotropic Fe0.5Mn0.5TiO3. At the other end of the spectrum, we have very isotropic alloys such as AgMn or CuMn. [^9] It has been emphasized that modeling AgMn or CuMn is notoriously difficult [309], because of the presence of short-range spin-density wave ordering [79, 80, 221]. The experimental finding $\gamma=4.0$ in the strongly uniaxial Fe0.5Mn0.5TiO3 [165], midway between the Ising Edwards-Anderson model and the AgMn values, namely $\gamma_{\text{IEA}}=6.33\pm 0.11$ and $\gamma_{\text{AgMn}}=2.3\pm 0.2$, gives weight to the idea that symmetries may play a fundamental role in clarifying this problem.
Indeed, in the context of universality and the renormalization group, simplified symmetry-motivated Hamiltonians are crucial to compute critical exponents. The numerical simulation of microscopically motivated interactions such as RKKY is simply too costly. Hence, the Edwards-Anderson model, recall Eq. (3), was introduced with $n$-component spins $\vec{\sigma}_{i}=(S_{1,i},\ldots,S_{n,i})$ subjected to the constraint $\vec{\sigma}_{i}\cdot\vec{\sigma}_{i}=1$ (with $n=3$ we have the Heisenberg case, $n=2$ is known as XY, while with $n=1$ one recovers the Ising spins). In this way, the distinctive features of the RKKY interaction (quenched disorder, frustration, a three-dimensional lattice structure and $n$-dimensional rotational and reflection symmetries) are preserved. This Heisenberg Edwards-Anderson (HEA) model was hoped to be a streamlined representation for highly isotropic materials, such as CuMn or AgMn, yet sufficient for computing critical exponents. Unfortunately, the story turns out not to be that simple.
All attempts to identify a phase transition at a finite temperature in the HEA model carried out during the 1980s and 1990s failed [255, 286, 265, 249]. However, Matsubara et al. [249] found a phase transition, provided that a small, random and anisotropic term is added to the HEA Hamiltonian. The crucial importance of the anisotropic interactions was also emphasized by Gingras [154]. The accepted picture at the time was that the lower critical dimension (i.e., the spatial dimension below which there is no phase transition) for the HEA model lay somewhere between 3D and 4D [99]. This theoretical picture, when confronted with the experimental finding of a finite-temperature transition in materials with only small anisotropies [227], suggests that anisotropic interactions (no matter how small) play a fundamental role.
A different, and provocative viewpoint was suggested by Mauger et al. [251]: although there was no spin-glass transition in the fully isotropic HEA model, a different order parameter called chirality might be present. Chirality is a scalar observable that describes vorticity and alignment between neighboring spins. This idea was elaborated upon by Kawamura [200] in his spin-chirality decoupling scenario. In the ideal case of a purely isotropic system, only chiralities would order at a nonvanishing temperature $T_{\text{CG}}>0$, but any small anisotropy would couple spins and chiralities (resulting in spins ordering at $T_{\text{SG}}=T_{\text{CG}}$). Yet, at the turn of the century, the scenario changed quite abruptly. Lee and Young [225] employed more efficient simulation algorithms to equilibrate larger systems at lower temperatures than was possible in the 1980s. By combining these new results with modern finite-size scaling techniques, they showed that, contrary to all previous expectations, $T_{\text{SG}}>0$ for the purely isotropic HEA model. The main open question today regards a comparison of $T_{\text{SG}}$ and $T_{\text{CG}}$ for the isotropic model. While some groups find that the most economic interpretation of their simulations for the HEA model is the absence of spin-chirality decoupling, i.e., $T_{\text{CG}}=T_{\text{SG}}$ [81, 129, 274], Kawamura and coworkers found that $T_{\text{CG}}$ is larger than —yet very close to— $T_{\text{SG}}$ [351, 284].
Real materials, however, always have some degree of anisotropy, hence one needs to understand what happens when some random (albeit small) anisotropic interactions are added to the HEA model. An analogy with ferromagnetic systems suggests that the anisotropic term will be a relevant perturbation in the renormalization group sense. This scenario has some consequences. One starts by modeling a real material through an HEA model, plus random anisotropies. We quantify these anisotropies through a coupling constant $g_{\text{micro}}$ when the correlation length is $\xi\sim 1$ lattice spacing ($g_{\text{micro}}$ will admittedly be very small for CuMn or AgMn). Now, even a tiny $g_{\text{micro}}$ gets amplified to $g_{\text{effective}}=g_{\text{micro}}\xi^{\phi}$ in an effective Hamiltonian for a correlation length $\xi$ $\phi$ is a positive crossover exponent, see, e.g., 16]. Hence, a new renormalization group fixed-point takes over and dictates the asymptotic critical exponents as $\xi\to\infty$. Presumably, the relevant fixed point at large-enough $\xi$ will be the one of the Ising Edwards-Anderson model. Two known facts strongly hint at the validity of this scenario: (i) $T_{\text{SG}}$ grows inordinately for even a tiny $g_{\text{micro}}$ (247; this could be guessed from 249), and (ii) for large-enough $g_{\text{micro}}$ the critical exponents turn out to be those of the Ising Edwards-Anderson model [294, 230, 39]. It is an unfortunate fact, however, that we know essentially nothing about the details of this crossover from Heisenberg to Ising behavior. Investigating the crossover experimentally will also be a serious challenge, because the dynamic critical exponent appropriate for $T>T_{\mathrm{g}}$ is large ($z\approx 6$, recall Fig. 24), which means that doubling $\xi$ —by lowering $T>T_{\mathrm{g}}$— while keeping the sample in thermal equilibrium requires extending the time duration of experiments by a factor of 64.
Kawamura and his collaborators have advanced a phenomenological interpretation of the existing experiments that somehow circumvents this frustrating lack of knowledge about the crossover that arises when $g_{\text{micro}}>0$. They start with simulations of the HEA model with $g_{\text{micro}}=0$ [203, 284]. According to their view, $T_{\text{CG}}>T_{\text{SG}}$ for the purely isotropic HEA model but they conjecture that, as soon as $g_{\text{micro}}>0$, there is a single phase transition (i.e., $T_{\text{CG}}=T_{\text{SG}})$. Although in typical experiments the magnetization (a spin-related quantity) is measured, the relevant critical exponents would be those of the chiralities (as computed for $g_{\text{micro}}=0$). In addition, Kawamura [201] proposed ways to access the chiralities that have been demonstrated experimentally [336, 312]. This approach was followed in Taniguchi [335] by measuring the exponent $\delta$ from the chiralities, finding results in agreement with determinations of $\delta$ from the magnetization. As for exponent $\gamma$, one has $\gamma_{SG}^{g_{\text{micro}}=0}\approx 3$ and $\gamma_{CG}^{g_{\text{micro}}=0}\approx 2$ [129, 284]. Of the two $\gamma$ exponents, both of them computed for the purely isotropic HEA model, the chiral one lies closer to the experimental determination for 0.5 at.% AgMn, $\gamma=2.3\pm 0.2$ [227]. More comparisons of critical exponents can be found in Kawamura [202], Kawamura and Taniguchi [203], Ogawa et al. [284].
There have been as well claims of different nonequilibrium behavior for Ising and Heisenberg materials in (at least) two different aspects:
  1. 1.
    Dupuis et al. [118] and Bert et al. [54] described the aging dynamics as an activated process with free-energy barriers $\sim\xi_{\mathrm{Zeeman}}^{\varPsi}$, with an exponent $\varPsi$ that would be different for Ising and Heisenberg materials. However, Bert et al. [54] extracted $\xi_{\mathrm{Zeeman}}$ in different ways for Heisenberg and Ising materials, which has been recently shown to be unnecessary [289]. Whether or not the overall dynamical differences in Heisenberg or Ising materials were simply caused by the different ways of extracting $\xi_{\mathrm{Zeeman}}$ is an open question.
  2. 2.
    Hérisson and Ocio [174] noted in their extended account of their experiment [173] that the measured fluctuation-dissipation ratio (FDR) —recall Sec. IV.3— was surprisingly curvature-less. This feature is reminiscent of the FDR predicted for systems with one-step RSB [the lack of curvature had been independently related to chiralities by 201]. We should note, however, that the Ising Edwards-Anderson model displays changes even in the sign of the curvature for its FDR obtained at finite times —see Fig. 29 and Sec. IV.4. It is, hence, clear that one can find measuring-time combinations in which the FDR is also curvature-less for an Ising system, a fact that calls for caution.
In our opinion, the important theoretical challenge for the future will be developing an appropriate field-theoretical framework that can explain the crossover when the random-anisotropy $g_{\text{micro}}$ becomes different from zero. One would like to understand as well how the chiralities mix with the spin operators. Another important open problem would be to understand why the Ising Edwards-Anderson model reproduces quantitatively the experimental aging dynamics at $T<T_{\mathrm{g}}$, while the same model does not make a good job of explaining the critical exponents. Perhaps the fact that critical exponents are equilibrium quantities (which enforces carrying out experiments at $T>T_{\mathrm{g}}$) will play a role in solving this conundrum. Clearly enough, further numerical investigations of the HEA model with random anisotropies can be instrumental in understanding the crossover. From the experimental side, new determination of the critical exponents on single crystals that allow reaching large values of $\xi$ could shed light on the crossover, particularly by focusing on changes of the effective values for the critical exponents as $\xi$ grows.
VII.1.4 Experiments with potential for strong impact on theory